OpenStudy (anonymous):

Wats the derivative of pi^2?

OpenStudy (anonymous):

I typed 2pi but its not working!!!!!!!!!!!!!?????????

OpenStudy (kainui):

Must be some kind of error in the back of the book, this is clearly the right answer. \[\lim_{h \rightarrow 0}\frac{ (\pi+h)^2 -h^2}{ h }=2\pi\] I just checked with the definition and everything.

OpenStudy (anonymous):

yeah ur rite thxxx!!!!!

OpenStudy (anonymous):

derivative of pi^2 is 2pi . correct

OpenStudy (anonymous):

Yeah it's definitely an error. 2pi is correct

OpenStudy (anonymous):

u guys r the best!!!! if u hav a kik i'll mesege u my name! ;)

OpenStudy (ranga):

Isn't the derivative of a constant zero?

OpenStudy (lncognlto):

The derivative of pi^2 would indeed be 2pi. :)

OpenStudy (anonymous):

oh r u sure ranga? thts not wat my prof said the other day

OpenStudy (lncognlto):

@Ranga, are you sure the rules of constants apply to Pi?

OpenStudy (superdavesuper):

wait wait wait @kainui Pi is a constant so its derivative should be zero right?

OpenStudy (superdavesuper):

I agree w @ranga on this one :)

OpenStudy (ranga):

pi is a number, an irrational number. It is not a variable. The derivative of a constant is zero.

OpenStudy (kainui):

@superdavesuper Incorrect. Since 2pi is periodic, just like the derivative of a constant of 0, it's exactly the same as if you go 2pi, 4pi, etc...

OpenStudy (anonymous):

I agree with @incognito @AustinC @Kainui The derivative of pi^2 would be 2pi

OpenStudy (anonymous):

Yeah like troy and them said pi^2 is definitely 2pi

OpenStudy (anonymous):

ty so much! 2pi for sure. I'll call my friends to tell them tha anser is rong in the bak of the book. :) :) :)

OpenStudy (anonymous):

\(\large \lim \limits_{h \rightarrow 0}\frac{ \pi^2 -\pi^2}{ h }= 0\)

OpenStudy (ranga):

what is the answer in the back of the book?

OpenStudy (anonymous):

@janaki.t23 u forgot the h in the top part thats y ur getin the rong answer.

OpenStudy (anonymous):